### v1

# Towards Quantum Universal Hypothesis Testing thanks: B. Chen would like to acknowledge the support of a Google gift fund

Abstract
Hoeffding’s formulation and solution to the universal hypothesis testing (UHT) problem had a profound impact on many subsequent works dealing with asymmetric hypotheses. In this work, we introduce a quantum universal hypothesis testing framework that serves as a quantum analog to Hoeffding’s UHT. Motivated by Hoeffding’s approach, which estimates the empirical distribution and uses it to construct the test statistic, we employ quantum state tomography to reconstruct the unknown state prior to forming the test statistic. Leveraging the concentration properties of quantum state tomography, we establish the exponential consistency of the proposed test: the type II error probability decays exponentially quickly, with the exponent determined by the trace distance between the true state and the nominal state.
Index Terms:
Quantum Hypothesis Testing, Universal Hypothesis Testing, Quantum Universal Hypothesis Testing, Quantum State Tomography.
I Introduction
Quantum hypothesis testing (QHT) dates back to 1967, when Helstrom formally introduced quantum detection theory for “choosing one of two density operators as the better description of the state" [1]. The projection detector proposed by Helstrom reduces to the classical likelihood ratio test when the two density matrices commute. Since then, many significant results have been discovered in QHT. Examples include Helstrom’s bound on minimum error probability in distinguishing two quantum states [2] and subsequent development in obtaining asymptotic performence bounds under different setups (see, e.g., [3]). Perhaps the most celebrated result is the Quantum Stein’s lemma [4], which - thanks to the strong converse established in [5] - provides a clean result that serves as a precise quantum analog of the classical Stein’s lemma [6].
This parallel - both in problem formulations and in solutions - between classical and QHT is quite apparent in many of the results in the literature. Besides the Quantum Stein’s lemma[5], we can use the Chernoff bound [7] when both types of errors are considered in a framework that differs from the Neyman-Pearson setup [8]. Indeed, this observation is what motivates the present work: although there is a line of work focused on it in classical hypothesis testing, the so-called universal hypothesis testing (UHT) has remained largely unexplored in the quantum setting. UHT addresses the problem of determining whether a sequence of samples is generated from a known nominal distribution. It is, in essence, equivalent to the goodness of fit testing [9, 10, 11, 12]. Hoeffding’s formulation for UHT [13], in which the unknown alternative is a fixed distribution in the probability simplex, leads to a surprising result: even though this alternative is completely unknown, the achievable performance (in terms of the type II error probability, or more precisely, its exponential decay rate) is essentially the same as that of simple hypothesis testing when the alternative is fully specified.
The generality of the UHT framework makes it particularly appealing for a broad class of problems encountered in both classical and quantum information systems in which the goal is to verify or authenticate the distribution/state when given multiple copies of samples/quantum systems. The present work defines the quantum universal hypothesis testing (QUHT) problem and provides a simple solution to this problem through the use of quantum state tomography (QST). Our approach is largely motivated by Hoeffding’s treatment of the classical UHT, which is, using modern terminology, a generalized likelihood ratio test (GLRT). With a GLRT, the empirical distribution is estimated using the sample sequence and used in constructing the likelihood ratio test that simplifies to the Kullback-Leibler divergence (KLD) between the empirical distribution and the nominal distribution. The concentration of the empirical distribution around the true distribution, as the number of samples increases, leads to the desired error probability performance.
While QUHT has not been formally defined or studied to date, there have been several related works reported in the quantum literature. These include extensions of Sanov’s theorem [14, 15, 16, 17], composite hypothesis testing [18], and Hoeffding-type error exponents in binary quantum testing [19, 20, 15]. Additionally, there are several lines of research in quantum circuits and information where QUHT can play an important role. Examples include quantum circuit equivalence [21, 22], quantum state discrimination [23, 24, 25, 26], and quantum state verification and certification [27, 28, 29].
The UHT formulation is often referred to as the one-sample test, especially in the machine learning (ML) community. One can similarly define the quantum analog to the two-sample test [30] in which two sequences of samples (quantum systems) are given and the task is to determine whether the two sequences are generated from the same distribution (quantum state). Our proposed solution to the UHT can be easily generalized to the two-sample setting, since respective concentrations of the two-sample sequences will ensure exponential consistency of the hypothesis test when QST is applied to both sequences.
Figure 1: QST ensures the concentration of the reconstructed state around the true state $\sigma$. When $\sigma\neq\rho$, separation of the reconstructed state from the nominal state $\rho$ is ensured asymptotically.
II Problem Setup
Let $\mathcal{H}$ be a finite-dimensional Hilbert space. We denote by $\mathcal{L}(\mathcal{H})$ the space of linear operators on $\mathcal{H}$ and by $\mathcal{U}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ the group of unitary operators satisfying $U^{\dagger}U=UU^{\dagger}=I_{\mathcal{H}}$. The set of density operators $\mathcal{D}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ consists of positive semidefinite, unit-trace operators [31]. In the special case $\dim(\mathcal{H}_{2})=2$, $\mathcal{D}(\mathcal{H}_{2})$ is the set of qubit density operators. We write $\operatorname{Tr}(\cdot)$ for the trace functional, and for any $\sigma\in\mathcal{D}(\mathcal{H})$, its $m$-fold tensor product is denoted $\sigma^{\otimes m}\in\mathcal{D}(\mathcal{H}^{\otimes m})$. Throughout, $\log$ and $\ln$ denote logarithms base 2 and $e$, respectively. The trace norm of $\rho\in\mathcal{L}(\mathcal{H})$ is $\|\rho\|_{1}=\operatorname{Tr}[\sqrt{\rho^{\dagger}\rho}]$, and the trace distance between $\rho$ and $\sigma$ is $T(\rho,\sigma)=\tfrac{1}{2}\|\rho-\sigma\|_{1}$. The quantum fidelity is $F(\rho,\sigma)=\|\sqrt{\rho}\,\sqrt{\sigma}\|_{1}^{2}$, and the quantum relative entropy is $D(\rho\|\sigma)=\operatorname{Tr}[\rho\ln\rho-\rho\ln\sigma]$ [31].
II-A Quantum Hypothesis Testing
Analogous to the classical setting, in QHT, independent copies of quantum systems are generated from the same quantum state $\sigma$, and the task is to discern if $\sigma=\rho_{0}$ or $\sigma=\rho_{1}$, where $\rho_{0}\neq\rho_{1}\subset\mathcal{D}(\mathcal{H})$. Or formally,
$H_{0}:\sigma$ $=\rho_{0}$ (1)
$H_{1}:\sigma$ $=\rho_{1}.$ (2)
Any test is specified by a decision rule $M_{m}:\mathcal{D}(\mathcal{H})^{m}\to\{0,1\}$, such that $\hat{H}=H_{0}$ if $M_{m}(\sigma^{\otimes m})=0$, and $\hat{H}=H_{1}$ if $M_{m}(\sigma^{\otimes m})=1$. Here we use $\hat{H}$ to denote the declared hypothesis and will reserve $H$ for the true hypothesis. With this notation, we can now define the type I and type II error probabilities as follows:
$\alpha(M_{m})$ $:=\mathbb{P}_{M_{m}}[\hat{H}=H_{1}|H=H_{0}]$ (3)
$\beta(M_{m})$ $:=\mathbb{P}_{M_{m}}[\hat{H}=H_{0}|H=H_{1}].$ (4)
When it is clear from context, we may denote them by $\mathbb{P}[H_{1}|H_{0}]$ and $\mathbb{P}[H_{0}|H_{1}]$, respectively. Here use the term symmetric to describe the setting where we attempt to minimize the sum of errors $\alpha(M_{m})+\beta(M_{m})$, and asymmetric to denote the setting where we restrict the type I error to be less than $\alpha$ and we minimize the type II error.
With a single (or finite) copy of the quantum systems, the Helstrom bound [32] [33] provides a lower bound on the sum of the two error probabilities (which is in essence the average error probability, if an equal prior is assumed):
$\alpha(M_{m})+\beta(M_{m})\geq\frac{1}{2}\left(1-\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}\right).$ (5)
The Quantum Stein’s lemma [4, 5], on the other hand, provides an asymptotic performance bound when a large number of copies are available. Specifically, subject to a constraint on the type I error probability, there exists a sequence of decision rules ${M_{m}}$ such that:
$\lim_{m\to\infty}\frac{1}{m}\ln(\beta_{\rho_{0},\rho_{1}}(M_{m}))=-D(\rho_{0}\|\rho_{1}),$ (6)
Additionally, no other test achieves a better exponential decay rate.
The Quantum Stein’s lemma provides an operational meaning for the quantum relative entropy that is a precise analog of the classical case.
II-B Quantum Universal Hypothesis Testing
QUHT is a natural extension of the classical universal hypothesis testing framework to the quantum setting. Here, the null hypothesis asserts that the quantum systems are specified by a known nominal state $\rho$, while the alternative states otherwise, i.e.,
$H_{0}:\sigma$ $=\rho$ (7)
$H_{1}:\sigma$ $\neq\rho.$ (8)
This formulation directly mirrors Hoeffding’s formulation of the classical UHT, where the alternative is represented by an unknown but fixed state $\sigma$ that is different from $\rho$. For a given decision rule $M_{m}$, we denote the type I and II error probabilities as $\alpha(M_{m})$ and $\beta_{\rho,\sigma}(M_{m})$, respectively. We note that while any decision rule should not depend on $\sigma$ as it is unknown, the actual performance of the rule (in particular, type II error probability) is typically a function of $\sigma$. We may occasionally use $\alpha$ to refer to a constraint on the type I error, as well as the function describing the type I error of a decision rule, with the distinction to be clarified by context.
The generality of the QUHT formulation lies in the generality of the alternative hypothesis. While this motivates its application to broad applications in quantum systems, it also presents unique challenges. Many techniques developed for QHT rely on knowledge of both states — e.g., optimal tests for symmetric or structured alternatives [34, 14, 16] - and are thus not directly applicable when the alternative is entirely unspecified. An exception is found in recent work [17], which constructs a "quantum empirical distribution" and proves large-deviation results reminiscent of those in the classical setting. We further discuss its connection to our work in Section IV.
II-C Quantum Two-Sample Test
The QUHT described in the previous subsection is often referred to as the one-sample test in the classical setting. Extending this formulation to the two-sample case is straightforward. Suppose now we have two sequences of quantum systems: $\rho_{1},\ldots,\rho_{m}$ are generated from an unknown density matrix $\rho$ while $\sigma_{1},\ldots,\sigma_{n}$ are generated from another unknown density matrix $\sigma$. The two-sample test is used to discriminate between the following two hypotheses:
$H_{0}:\sigma$ $=\rho$ (9)
$H_{1}:\sigma$ $\neq\rho.$ (10)
While the form appears identical to the one-sample test (i.e., QUHT), the setup differs in one key way. The two-sample test has no knowledge of both $\rho$ and $\sigma$ and is given two sequences of samples, whereas the one-sample test has perfect knowledge of $\rho$ and is given a single sequence generated from the unknown state $\sigma$.
III Results
This section presents our main results on QUHT for both the one- and two-sample settings. As alluded to in Section I, we primarily leverage the concentration results found in the QST literature to obtain the error exponent for type II error probabilities. The only exception is when the nominal state is pure, in which case we obtain an optimal error exponent.
Table I presents our main results for both one-sample and two-sample variants of QUHT. To emphasize the dominant error-exponent terms, we have omitted those that vanish asymptotically as the number of copies of the quantum states under test grows without bound. Here, $m$ and $n$ denote the numbers of identical copies of the states $\rho$ and $\sigma$, respectively, as introduced in Section II-C. Throughout our analysis, the type I error threshold $\alpha\in(0,1)$ is treated as a fixed constant. In the two-sample setting, we restrict attention to the regime in which $m$ and $n$ increase at the same rate and define $k=\min\{m,n\}$ to express all bounds in terms of the effective sample size.
TABLE I: Table of QUHT Results. (1S: one-sample, 2S: two-sample)
Setting Measurement Type Tomography Source Type II Error
1S Qudit Pure State
Theorem 1
Indep. $[F(\rho,\sigma)]^{m}$
1S Qubit General State
Theorem 2
Indep. Lemma 1 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{54}\right\}$
2S Qubit General State
Theorem 3
Indep. Lemma 1 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{216}\right\}$
1S Qudit General State
Theorem 4
Indep. Lemma 2 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{86\cdot d^{3}}\right\}$
2S Qudit General State
Theorem 5
Indep. Lemma 2 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344\cdot d^{3}}\right\}$
1S Qudit General State
Theorem 6
Entangled Lemma 3 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}\right\}$
2S Qudit General State
Theorem 7
Entangled Lemma 3 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}\right\}$
III-A Pure State QUHT
We begin by establishing that for symmetric QUHT, we can achieve asymptotically optimal error performance by performing measurements in the eigenbasis of $\rho$—a result that closely parallels known optimality results for pure state symmetric QHT [35] [36].
Theorem 1 (Symmetric Pure State).
Take $\rho=\lvert\phi\rangle\langle\phi\rvert\in\mathcal{D}(\mathcal{H})$ a pure state density operator and $\sigma\in\mathcal{D}(\mathcal{H})$ a (potentially mixed) density operator. In the QUHT setting, given $\sigma^{\otimes m}$ and $\alpha\in(0,1)$, using independent measurements there exists a decision rule $M_{m}$, such that
$\alpha(M_{m})$ $=0$ (11)
$\beta_{\rho,\sigma}(M_{m})$ $=\left[F(\rho,\sigma)\right]^{m}.$ (12)
For such a test, we have
$\lim_{m\to\infty}\frac{1}{m}\ln\left(\alpha(M_{m})+\beta_{\rho,\sigma}(M_{m})\right)=-\tilde{D}_{1/2}(\rho\|\sigma),$ (13)
where $\tilde{D}_{1/2}(\rho\|\sigma):=-\ln(F(\rho,\sigma))$ is the sandwiched Rényi Relative Entropy [37], [38] between $\rho$ and $\sigma$ of order $1/2$.
In Section A-A we prove Theorem 1 and in Section A-B we replicate a result from [36] which demonstrates our result is tight in the sense that for any other decision rule $\tilde{M}_{m}$ we will have
$\liminf\limits_{m\to\infty}\frac{1}{m}\ln\left(\alpha(\tilde{M}_{m})+\beta_{\rho,\sigma}(\tilde{M}_{m})\right)\geq-\tilde{D}_{1/2}(\rho\|\sigma).$ (14)
III-B Qubit QUHT
Next, we provide an example illustrating how leveraging a concentration result for a quantum state estimator allows us to derive a corresponding result in the QUHT framework. Let $I_{2}$ be the $2\times 2$ identity and $\sigma_{X}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $\sigma_{Y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$, $\sigma_{Z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ the Pauli matrices, and matrix $\vec{\sigma}=[\sigma_{X},\sigma_{Y},\sigma_{Z}]^{T}$. Any Hermitian $A\in L(\mathcal{H}_{2})$ admits the Bloch decomposition
$A=\frac{1}{2}\bigl{(}I_{2}+\vec{r}\!\cdot\!\vec{\sigma}\bigr{)}$ (15)
Where $\vec{r}=[\operatorname{Tr}[A\sigma_{X}],\operatorname{Tr}[A\sigma_{Y}],\operatorname{Tr}[A\sigma_{Z}]]$
[…]
Let $\sigma\in\mathcal{D}(\mathcal{H}_{2})$ and $m$ be divisible by three. Denote the joint state of $m$ copies by $\sigma^{\otimes m}$, and partition its factors into three sets $G_{i}$ ($i\in\{X,Y,Z\}$) of size $m/3$. On each $G_{i}$, perform the $0,1$-valued projective measurement $M_{i}$ in the eigen-basis of the Pauli operator $\sigma_{i}\in\{\sigma_{X},\sigma_{Y},\sigma_{Z}\}$. The empirical Bloch coordinate is $\hat{r}_{i}=\frac{3}{m}\sum_{\ell\in G_{i}}M_{i}(\sigma)$, which satisfies $\mathbb{E}[\hat{r}_{i}]=\operatorname{Tr}[\sigma_{i}\sigma]$. The reconstructed state is
$\hat{\sigma}=\tfrac{1}{2}\Bigl{(}I+\sum_{i\in\{x,y,z\}}\sigma_{i}\,\hat{r}_{i}\Bigr{)}.$ (16)
Then we have the following concentration result:
Lemma 1 (Pauli Operator Concentration).
For $\sigma$, $\hat{\sigma}$ as described above, and for $\varepsilon>0$, we have
$\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}>\varepsilon\right]\leq 6\exp\left\{-\frac{m\varepsilon^{2}}{54}\right\}.$ (17)
As the number of samples increases, the confidence region around $\sigma$ becomes more concentrated, with $\rho$ eventually residing outside it, as seen in Figure 1. The proof of this result can be found in Section A-C.
Using this result, we can recover a bound on universal hypothesis testing for (mixed) qubit systems.
Theorem 2 (Qubit One-Sample).
In the QUHT one-sample setting, for $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{2})$, given $\sigma^{\otimes m}$ and $\alpha\in(0,1)$, using only independent measurements there exists a decision rule $M_{m}$ so that for $m$ large enough, we will have
$\alpha(M_{m})$ $\leq\alpha$ (18)
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{\frac{-m\left\|\rho-\sigma\right\|_{1}^{2}}{54}\right\},$ (19)
where the $\lesssim$ hides an additive $O\left(\left\|\rho-\sigma\right\|_{1}m^{1/2}\log^{1/2}(1/\alpha)\right)$ term in the exponent.
The proof can be found in Section A-D.
We can even use this concentration result to define a strategy for (mixed) two-sample qubit QUHT.
Theorem 3 (Qubit Two-Sample).
In the QUHT two-sample setting for $\rho,\sigma\in\mathcal{D}(\mathcal{H}_{2})$, suppose $\sigma^{\otimes m}$, $\rho^{\otimes n}$, $\alpha\in(0,1)$ were given, and take $k=\min\{m,n\}$. For large enough $m,n$, and using independent measurements, there exists a decision rule $M_{m}$ such that
$\alpha(M_{k})$ $\leq\alpha$ (20)
$\beta_{\rho,\sigma}(M_{k})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{216}\right\},$ (21)
where $\lesssim$ is hides an additive $O(k^{1/2}\ln(1/\alpha))$ term in the exponent.
The proof can be found in Section A-E.
III-C Qudit Independent Measurements
Unfortunately, extending the above result to higher-dimensional Hilbert spaces gives an error exponent that scales poorly with the dimension of the space. For example, any Hermitian operator on $b$ qubits ($d=2^{b}$) can be decomposed using $d^{2}-1$ components, the Pauli strings [39]. Then applying the same method would give
$\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\varepsilon\right]\leq 2(d^{2}-1)\exp\left\{-\frac{m\varepsilon^{2}}{2(d^{2}-1)^{3}}\right\}.$ (22)
However, for general $d$-dimensional qudit states, we can do better using independent measurements by applying the following result from quantum state tomography [40]:
Lemma 2 (Theorem 1  [40]).
For a density operator $\tau\in\mathcal{D}(\mathcal{H})$, with $\dim(\mathcal{H})=d$, and $\operatorname{rank}(\tau)=r$, it is possible to construct, via $m$ independent measurements, $\hat{\tau}\in\mathcal{D}(\mathcal{H})$ such that, for any $\delta\in(0,1)$,
$\mathbb{P}\left[\left\|\tau-\hat{\tau}\right\|_{1}\geq\delta\right]\leq d\exp\left\{-\frac{m\delta^{2}}{86dr^{2}}\right\}.$ (23)
Note that this is a special case of Guţă’s [40] more general result, where they take measurements that are $2$-designs [40, Definition 1]. Additionally, since $r\leq d$, we can weaken the bound to the following:
$\mathbb{P}\left[\left\|\tau-\hat{\tau}\right\|_{1}\geq\delta\right]\leq d\exp\left\{-\frac{m\delta^{2}}{86\cdot d^{3}}\right\}.$ (24)
This yields a concentration result that we can use in the same manner as before.
Theorem 4 (Independent One-Sample).
In the QUHT one-sample setting with $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, given $\sigma^{\otimes m}$ and $\alpha\in(0,1)$, using only independent measurements there exists a decision rule $M_{m}$, so that for $m$ large enough, we will have
$\alpha(M_{m})$ $\leq\alpha$ (25)
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{86d^{3}}\right\},$ (26)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\log(1/\alpha)\right)$ term in the exponent.
For proof see Section A-F. This also yields a result for the two-sample QUHT setting.
Theorem 5 (Independent Two-Sample).
In the QUHT two-sample setting with $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, suppose $\sigma^{\otimes m}$, $\rho^{\otimes n}$, $\alpha\in(0,1)$ were given, and take $k=\min\{m,n\}$. For large enough $m,n$, and using independent measurements, there exists a decision rule $M_{m}$ such that
$\alpha(M_{k})$ $\leq\alpha$ (27)
$\beta_{\rho,\sigma}(M_{k})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344d^{3}}\right\},$ (28)
where $\lesssim$ is hiding an additive $O(k^{1/2}\ln(1/\alpha))$ term in the exponent.
See Section A-G for proof.
III-D Entangled Measurements
While independent measurements are easier to implement, it has been shown that using entangled measurements allows for quantum state tomography with a sample complexity that is unachievable using independent measurements [41].
Lemma 3 (Section 2.3 [41]).
For a density operator $\tau\in\mathcal{D}(\mathcal{H})$, with $\dim(\mathcal{H})=d$, and $\operatorname{rank}(\tau)=r$, it is possible to construct a density operator $\hat{\tau}\in\mathcal{D}(\mathcal{H})$ using $\tau^{\otimes m}$, such that for all $\delta\in(0,1)$,
$\mathbb{P}\left[F(\tau,\hat{\tau})\leq 1-\delta\right]\leq(m+1)^{3rd}\exp\left\{-2m\delta\right\}.$ (29)
We further weaken this bound to the following:
$\mathbb{P}\left[F(\tau,\hat{\tau})\leq 1-\delta\right]\leq(m+1)^{3d^{2}}\exp\left\{-2m\delta\right\}.$ (30)
This concentration result allows us to establish tighter bounds on the type II error rate. Specifically, we have the following one-sample result:
Theorem 6 (Entangled One-Sample).
In the QUHT one-sample setting with $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, given $\sigma^{\otimes m}$, $\alpha\in(0,1)$, and allowing measurements involving entanglement, there exists a decision rule $M_{m}$ such that for $m$ large enough,
$\alpha(M_{m})$ $\leq\alpha$ (31)
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}\right\},$ (32)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\log^{1/2}(m)\right)$ term in the exponent.
See Section A-H for the proof. We can also use this to get results for two-sample universal hypothesis testing. In particular, using entangled measurements, we can achieve the following:
Theorem 7 (Entangled Two-Sample).
In the QUHT two-sample setting with $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, suppose we are given $\sigma^{\otimes m}$, $\rho^{\otimes n}$, $\alpha\in(0,1)$, and take $k=\min\{m,n\}$. For large enough $m,n$, and using entangled measurements, there exists a decision rule $M_{m}$ such that
$\alpha(M_{m})$ $\leq\alpha$ (33)
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}\right\},$ (34)
where $\lesssim$ hides an additive $O(k^{1/2}\ln^{2}(k))$ term in the exponent.
See Section A-I for the proof.
IV Discussion
Recall that in the classical universal setting, Hoeffding was able to characterize the error exponent in terms of the KLD. In this work, we have established positive error exponents in terms of not the KLD but the trace norm. Specifically, we showed various results of the following form:
$\lim_{n\to\infty}\left\{\frac{1}{m}\ln(\beta_{\rho,\sigma}(M_{m}))\right\}=-m\frac{\left\|\rho-\sigma\right\|_{1}^{2}}{c},$ (35)
where $c\in\mathbb{R}$ depends on the dimension of the underlying Hilbert space, and the types of measurements used. This general form applies to both the one-sample and two-sample settings, with the latter offering greater flexibility in scenarios where the nominal state is itself inaccessible or generated from experimental data.
Comparing with the classical results for the universal hypothesis testing problem, which achieves an error exponent scaling as the relative entropy between $\rho$ and $\sigma$, this result appears somewhat lacking. For example, we have quantum Pinsker’s inequality [42, Theorem 3.1], which establishes that the squared trace distance forms a lower bound for the relative entropy. Since Hoeffding establishes his result in the classical setting using a large deviation type bound, a first attempt at a quantum generalization might utilize quantum variations of the classical Sanov’s theorem [6, Theorem 11.4.1]. However, in [14] and [16], the result required knowledge of the alternative hypothesis, which is not reasonable for QUHT.
As a corollary to [17, Lemma 1, Corollary 6], one could get a decision rule $\hat{M}_{m}$ such that
$\alpha(\hat{M}_{m})$ $\leq\alpha$ (36)
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln(\beta_{\rho,\sigma}(\hat{M}_{m}))\right\}$ $\leq-\hat{D}(\rho,\sigma),$ (37)
where $\hat{D}(\rho\|\sigma)$ is a cousin of the relative entropy, defined as
$\hat{D}(\rho\|\sigma)=\lim_{s\to 0}\left\{\frac{-\ln\left(\operatorname{Tr}\left[\sigma^{\frac{1-s}{2s}}\rho\sigma^{\frac{1-s}{2s}}\right]^{s}\right)}{s}\right\}.$ (38)
We briefly show the result in Section A-J. However, [17, Lemma 3] showed that $\hat{D}(\rho\|\sigma)\leq D(\rho\|\sigma)$, and the relationship between $\hat{D}(\rho\|\sigma)$ and the trace norm is not known, which makes direct comparison difficult. Nevertheless, the quantum empirical distribution they define is likely worth consideration for establishing tighter bounds for QUHT.
If we could find a decision rule $M_{m}^{*}$ with
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln(\beta_{\rho,\sigma}(M_{m}^{*}))\right\}=-D(\rho\|\sigma),$ (39)
then applying Quantum Pinsker’s Inequality, which is tight up to a constant factor [^1] When $\rho,\sigma$ share an eigen basis, the inequality reduces to the classical version, which is known to be tight to the constant [43], we would have $\beta_{\rho,\sigma}(M^{*}_{m})\simeq e^{-mD(\rho,\sigma)}\leq e^{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}}$. However, in Theorem 6, we showed that there exists a decision rule $M_{m}$ such that $\beta_{\rho,\sigma}(M_{m})\simeq e^{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}}$. In other words, even if a decision rule achieving an error exponent of $D(\rho\|\sigma)$ were available, its implied performance in terms of trace distance would be upper-bounded by a term matching the exponent achieved using tomography.
Several other works have considered similar problems under different names in other areas of quantum information science [21] [22] [27] [28], often motivated by practical needs for verifying quantum devices or certifying quantum behavior in minimally structured settings.. By framing QUHT as a formal hypothesis testing problem, we have aimed to provide both conceptual clarity and a foundation for future work. While our focus has been on leveraging quantum state tomography to achieve provable error exponents, the broader framework of one- and two-sample QUHT may have applications analogous to those of classical universal hypothesis testing. We view the present results as a first step in this direction, and hope they stimulate further theoretical and applied developments.
Appendix A Proofs
We will make heavy use of the following relationship between the trace distance and the fidelity of a quantum state [31]
Lemma 4.
$\rho,\sigma\in\mathcal{D}(\mathcal{H})$, then
$\frac{1}{2}\left\|\rho-\sigma\right\|_{1}\leq\sqrt{1-F(\rho,\sigma)}.$ (40)
We also have the following quantum version of Pinsker’s inequality [42, Theorem 3.1]
Lemma 5.
$D(\rho\|\sigma)\geq\frac{1}{2}\left\|\rho-\sigma\right\|_{1}^{2}.$
A-A Proof of Theorem 1
[…]
Now for $\tau\in\mathcal{D}(\mathcal{H})$, define $M_{\lvert 0\rangle}(\tau)$ to be the result of measuring $\tau$ using the positive operator valued measure (POVM) $\{\lvert 0\rangle\langle 0\rvert,I-\lvert 0\rangle\langle 0\rvert\}$, such that $M_{\lvert 0\rangle}(\tau)=1$ if the outcome is $\lvert 0\rangle$, and 0 otherwise. For $i\in[m]$, take $\sigma_{i}$ to be the $i^{\text{th}}$ copy of $\sigma$ in $\sigma^{\otimes m}$. Then our decision rule would be
$\hat{H}_{0}:$ $\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})=m$
$\hat{H}_{1}:$ $\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})<m.$
We naturally have $\alpha(M_{m})=0$. And by the Born rule, we have
$\mathbb{P}\left[M_{\lvert 0\rangle}(\sigma_{i})=1\right]=F(\rho,\sigma),$
so that for the type II error, by the independence of the measurements
$\beta_{\rho,\sigma}(M_{m})$ $=\mathbb{P}\left[\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})=m\right]$
$=\left[\prod_{i=1}^{m}\mathbb{P}\left[M_{\lvert 0\rangle}(\sigma_{i})=1\right]\right]$
$=\left[F(\rho,\sigma)\right]^{m}.$
A-B Tightness of Symmetric Hypothesis Testing Error Bound
The following result is essentially what was shown in Appendix F of [36].
From [32] [33] we know that the sum of type I and II errors is bounded by $\varepsilon^{*}_{m}:=\frac{1}{2}\left(1-\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}\right)$. Then consider the following argument:
$\varepsilon^{*}_{m}$ $=\frac{1}{2}\left(1-\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}\right)$ (41)
$=\frac{1}{2}\cdot\frac{(1-\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1})(1+\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1})}{1+\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}}$ (42)
$=\frac{1}{2}\cdot\frac{1-\frac{1}{4}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}^{2}}{1+\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}}$ (43)
$\geq\frac{1}{2}\cdot\frac{1-\frac{1}{4}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}^{2}}{2}$ (44)
$\geq\frac{1}{4}F(\rho^{\otimes m},\sigma^{\otimes m})$ (45)
$=\frac{1}{4}\left[F(\rho,\sigma)\right]^{m},$ (46)
where line 45 comes from Lemma 4.
A-C Proof of Lemma 1
Proof.
For all $K\in\{X,Y,Z\}$, suppose that
$\mathbb{P}\left[\frac{1}{2}\left\|\sigma_{K}-\hat{\sigma}_{K}\right\|_{1}>\varepsilon\right]\leq 2\exp\left\{-\frac{m\varepsilon^{2}}{6}\right\};$ (47)
then we have
$\left\|\sigma-\hat{\sigma}\right\|_{1}$ $=\frac{1}{2}\left\|(\sigma_{X}-\hat{\sigma}_{X})+(\sigma_{Y}-\hat{\sigma}_{Y})+(\sigma_{Z}-\hat{\sigma}_{Z})\right\|_{1}$
$\leq\frac{1}{2}\left\|\sigma_{X}-\hat{\sigma}_{X}\right\|_{1}+\frac{1}{2}\left\|\sigma_{Y}-\hat{\sigma}_{Y}\right\|_{1}+\frac{1}{2}\left\|\sigma_{Z}+\hat{\sigma}_{Z}\right\|_{1},$
Therefore
$\implies\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\varepsilon\right]$
$\leq\mathbb{P}\left[\frac{1}{2}(\left\|\sigma_{X}-\hat{\sigma}_{X}\right\|_{1}+\left\|\sigma_{Y}-\hat{\sigma}_{Y}\right\|_{1}+\left\|\sigma_{Z}-\hat{\sigma}_{Z}\right\|_{1})\geq\varepsilon\right]$
$\leq\mathbb{P}\left[\frac{1}{2}\left\|\sigma_{X}-\hat{\sigma}_{X}\right\|_{1}\geq\frac{\varepsilon}{3}\right]+\mathbb{P}\left[\frac{1}{2}\left\|\sigma_{Y}-\hat{\sigma}_{Y}\right\|_{1}\geq\frac{\varepsilon}{3}\right]$
$\ \ \ +\mathbb{P}\left[\frac{1}{2}\left\|\sigma_{Z}-\hat{\sigma}_{Z}\right\|_{1}\geq\frac{\varepsilon}{3}\right]$
$\overset{(a)}{\leq}3\left(2\exp\left\{-\frac{m\varepsilon^{2}}{54}\right\}\right)$
$=6\exp\left\{-\frac{m\varepsilon^{2}}{54}\right\},$
where $(a)$ comes from the application of Equation 47.
Now, all that remains is to prove Equation 47. Note that
$\frac{1}{2}\left\|\sigma_{K}-\hat{\sigma}_{K}\right\|_{1}$ $=\frac{1}{2}\left\|K\operatorname{Tr}[K\sigma]-K\hat{p}_{K}(\sigma)\right\|_{1}$
$=\frac{1}{2}\left|\operatorname{Tr}[K\sigma]-\hat{r}_{K}(\sigma)\right|\cdot\left\|K\right\|_{1}$
$=\left|\operatorname{Tr}[K\sigma]-\hat{r}_{K}(\sigma)\right|$
$=\left|\mathbb{E}[\hat{r}_{K}(\sigma)]-\hat{r}_{K}(\sigma)\right|.$
So
$\mathbb{P}\left[\frac{1}{2}\left\|\sigma_{K}-\hat{\sigma}_{K}\right\|_{1}\geq\varepsilon\right]$ $=\mathbb{P}\left[\left|\mathbb{E}[p_{K}(\sigma)]-\hat{p}_{K}(\sigma)\right|\right]$
$\leq 2\exp\left\{-\frac{m\varepsilon^{2}}{6}\right\}$
where the last inequality is an application of Hoeffding’s Inequality [44]. ∎
A-D Proof of Theorem 2
Suppose we constructed $\hat{\sigma}$ using measurements in the Pauli basis, as in Lemma 1. Suppose we use the following decision rule
$\hat{H}_{0}:\left\|\rho-\hat{\sigma}\right\|_{1}$ $\leq c_{m}$
$\hat{H}_{1}:\left\|\rho-\hat{\sigma}\right\|_{1}$ $>c_{m}.$
Using Lemma 1, we can directly bound the Type I error
$\mathbb{P}[\hat{H}=H_{1}|H=H_{0}]$ $=\mathbb{P}[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq c_{m}]$
$\leq 6\exp\left\{-\frac{mc_{m}^{2}}{54}\right\}.$
As for the type II error, note that if $H_{1}$ is true, then $\left\|\rho-\sigma\right\|_{1}=c>0$. Then so long as $c_{m}$ is a decreasing function of $m$, then $\exists$ some value of $m$ such that the requirements of Lemma 1 are met for $\varepsilon=\left\|\rho-\sigma\right\|_{1}-c_{m}$. Then we consider the following:
$\{\hat{\sigma}:\left\|\rho-\hat{\sigma}\right\|_{1}>c_{m}\}\supset\{\hat{\sigma}:\left\|\rho-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}>c_{m}\}$
Therefore
$\mathbb{P}\left[\hat{H}=H_{1}|H=H_{1}\right]$
$=\mathbb{P}\left[\left\|\rho-\hat{\sigma}\right\|_{1}>c_{m}\right]$
$\geq\mathbb{P}\left[\left\|\rho-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}>c_{m}\right]$
$=\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\leq\left\|\rho-\sigma\right\|_{1}-c_{m}\right]$
$\geq 1-6\exp\left\{-\frac{m(\left\|\rho-\sigma\right\|_{1}-c_{m})^{2}}{54}\right\}$
$\geq 1-6\exp\left\{\frac{-m\left\|\rho-\sigma\right\|_{1}^{2}+2mc_{m}\left\|\rho-\sigma\right\|_{1}}{54}\right\}.$
Now, since we require $\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]\leq\alpha$. Then we can pick $c_{m}=\sqrt{\frac{54\log(6/\alpha)}{m}}$, so that
$\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]\leq 6\exp\left\{-\frac{mc_{m}^{2}}{54}\right\}=\alpha,$
and gives us
$\mathbb{P}\left[\hat{H}=H_{0}|H=H_{1}\right]$
$\leq 6\exp\left\{\frac{-m\left\|\rho-\sigma\right\|_{1}^{2}}{54}+O\left(m^{1/2}\log^{1/2}(1/\alpha)\right)\right\}.$
A-E Proof of Theorem 3
Suppose we constructed $\hat{\sigma},\hat{\rho}$ using measurements in the Pauli basis, as in Lemma 1. If we use the following as our decision rule $M_{k}$
$\hat{H}_{0}:\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}$ $\leq c_{m,n}$
$\hat{H}_{1}:\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}$ $>c_{m,n}.$
Where $c_{m,n}$ will be selected to control the Type I error. Let $\mathbb{P}_{H_{0}}$ be the probability of an event conditioned on $H_{0}$ being true. We bound the type I error like this:
$\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]$
$=\mathbb{P}_{H_{0}}[\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}>c_{m,n}]$
$\leq\mathbb{P}_{H_{0}}[\left\|\hat{\rho}-\rho\right\|_{1}+\left\|\hat{\sigma}-\rho\right\|_{1}\geq c_{m,n}]$
$=\mathbb{P}_{H_{0}}[\left\|\hat{\rho}-\rho\right\|_{1}+\left\|\hat{\sigma}-\sigma\right\|_{1}\geq c_{m,n}]$
$\leq\mathbb{P}_{H_{0}}\left[\left\|\rho-\hat{\rho}\right\|_{1}\geq\frac{c_{m,n}}{2}\right]+\mathbb{P}_{H_{0}}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\frac{c_{m,n}}{2}\right]$
$\leq 12\exp\left\{-\frac{kc_{m,n}^{2}}{216}\right\}$
Then taking $c_{m,n}=\Theta\left(k^{-1/2}\ln\left(1/\alpha\right)\right)$, gives us $\alpha(M_{m})\leq\alpha$.
And in terms of the type II error, take $\mathbb{P}_{H_{1}}$ to be the probability of an event conditioned on $H_{1}$ being the true hypothesis. Now, consider the following argument.
By the triangle inequality, we have
$\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}$ $\geq\left\|\rho-\hat{\sigma}\right\|_{1}-\left\|\rho-\hat{\rho}\right\|_{1}$
$\geq\left\|\rho-\sigma\right\|_{1}-\left\|\rho-\hat{\rho}\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}$
Therefore
$\mathbb{P}\left[\hat{H}=H_{0}|H=H_{1}\right]$
$=\mathbb{P}_{H_{1}}[\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}\leq c_{m,n}]$
$\leq\mathbb{P}_{H_{1}}[\left\|\rho-\sigma\right\|_{1}-\left\|\rho-\hat{\rho}\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}\leq c_{m,n}]$
$=\mathbb{P}_{H_{1}}[\left\|\sigma-\hat{\sigma}\right\|_{1}+\left\|\rho-\hat{\rho}\right\|_{1}\geq\left\|\rho-\sigma\right\|_{1}-c_{m,n}]$
$\leq\mathbb{P}_{H_{1}}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{2}\right]$
$\ \ \ +\mathbb{P}_{H_{1}}\left[\left\|\rho-\hat{\rho}\right\|_{1}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{2}\right]$
And under $H_{1}$, we have $\left\|\rho-\sigma\right\|_{1}>0$. Then since $c_{m,n}\searrow 0$ as $k\to\infty$, then for $k$ large enough we have $\left\|\rho-\sigma\right\|_{1}-c_{m,n}>0$. Then for $k$ large enough, we can apply Lemma 1, to get
$\beta_{\rho,\sigma}(M_{k})$ $\leq 12\exp\left\{-\frac{k(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{216}\right\}$
$\leq 12\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{216}+\frac{m\cdot c_{m,n}}{108}\right\}$
$=\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{216}+O\left(k^{1/2}\ln(1/\alpha)\right)\right\}$
A-F Proof of Theorem 4
Here, we make use of Equation 24, to get $\hat{\sigma}$. Then our decision rule will be
$\hat{H}_{0}:\left\|\rho-\hat{\sigma}\right\|_{1}$ $<c_{m}$
$\hat{H}_{1}:\left\|\rho-\hat{\sigma}\right\|_{1}$ $\geq c_{m},$
where we will specify $c_{m}$ such that $\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]\leq\alpha$.
Now,
$\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]$ $=\mathbb{P}\left[\left\|\rho-\hat{\sigma}\right\|_{1}>c_{m}|\rho=\sigma\right]$
$=\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}>c_{m}\right]$
$\leq\exp\left\{-\frac{mc_{m}^{2}}{86d^{3}}+\ln(d)\right\}.$
Then if we take $c_{m}=\Theta\left(m^{-1/2}\ln(1/\alpha)\right)$, we get $\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]\leq\alpha$.
Now, under $H_{1}$ assume that $\left\|\rho-\sigma\right\|_{1}=c>0$,
$\mathbb{P}\left[\hat{H}=H_{0}|H=H_{1}\right]=\mathbb{P}\left[\left\|\rho-\hat{\sigma}\right\|_{1}\leq c_{m}|\left\|\rho-\sigma\right\|_{1}=c\right],$
From here on out we use $\mathbb{P}_{H_{1}}$ to denote the conditional event $\left\|\rho-\sigma\right\|_{1}=c$.
But
$\left\{\hat{\sigma}:\left\|\rho-\hat{\sigma}\right\|_{1}\leq c_{m}\right\}\subseteq\left\{\hat{\sigma}:\left\|\rho-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}\leq c_{m}\right\}$
=$=\left\{\hat{\sigma}:\left\|\sigma-\hat{\sigma}\right\|_{1}\leq\left\|\rho-\sigma\right\|_{1}-c_{m}\right\}$
$\implies\mathbb{P}_{H_{1}}\left[\left\|\rho-\hat{\sigma}\right\|_{1}\leq c_{m}\right]\leq\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\left\|\rho-\sigma\right\|_{1}-c_{m}\right].$
And since $c_{m}\searrow 0$ as $m\to\infty$, then if $m$ is large enough such that $c-c_{m}>0$, then
$\mathbb{P}\left[\left\|\rho-\hat{\sigma}\right\|_{1}\leq c_{m}|H=H_{1}\right]\leq\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq c-c_{m}\right]$
$\leq\exp\left\{-\frac{m(c-c_{m})^{2}}{86d^{3}}+\ln(d)\right\}$
$\leq\exp\left\{-\frac{mc^{2}}{86d^{3}}+\frac{mc_{m}}{43d^{3}}+\ln(d)\right\}$
$=\exp\left\{-\frac{mc^{2}}{86d^{3}}+O\left(m^{1/2}\ln(1/\alpha)\right)\right\}.$
A-G Proof of Theorem 5
We use the algorithm of [40] to construct both $\hat{\rho},\hat{\sigma}$ with the guarantee provided in Equation 24.
We let our decision rule be
$\hat{H}_{0}:$ $\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}<c_{m,n}$
$\hat{H}_{1}:$ $\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}\geq c_{m,n}.$
Of course, under $H_{0}$, we have
$\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}$ $\leq\left\|\hat{\rho}-\rho\right\|_{1}+\left\|\rho-\hat{\sigma}\right\|_{1}$
$=\left\|\hat{\rho}-\rho\right\|_{1}+\left\|\hat{\sigma}-\sigma\right\|_{1},$
so that
$\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]$
$=\mathbb{P}\left[\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}\geq c_{m,n}|\rho=\sigma\right]$
$\leq\mathbb{P}\left[\left\|\rho-\hat{\rho}\right\|_{1}+\left\|\sigma-\hat{\sigma}\right\|_{1}\geq c_{m,n}\right]$
$\leq\mathbb{P}\left[\left\|\rho-\hat{\rho}\right\|_{1}\geq\frac{c_{m,n}}{2}\right]+\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\frac{c_{m,n}}{2}\right]$
$\leq d\exp\left\{-\frac{mc_{m,n}^{2}}{344d^{3}}\right\}+d\exp\left\{-\frac{nc_{m,n}^{2}}{344d^{3}}\right\}$
$\leq\exp\left\{-\frac{kc_{m,n}^{2}}{344d^{3}}+\ln(2d)\right\}.$
Therefore, taking $c_{m,n}=\Theta\left(m^{-1/2}\log(1/\alpha)\right)$ we get  $\alpha=\mathbb{P}\left[\hat{H}=H_{1}|H=H_{0}\right]\leq\alpha$.
In the other direction, we will again assume that $\left\|\rho-\sigma\right\|_{1}=c>0$, and denote the probability of an event conditioned on this outcome by $\mathbb{P}_{H_{1}}\left[\cdot\right]$.
Now, we have
$\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}$ $\geq\left\|\rho-\hat{\sigma}\right\|_{1}-\left\|\rho-\hat{\sigma}\right\|_{1}$
$\geq\left\|\rho-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}-\left\|\rho-\hat{\rho}\right\|_{1},$
so that
$\beta_{\rho,\sigma}$ $=\mathbb{P}\left[\hat{H}=H_{0}|H=H_{1}\right]$
$=\mathbb{P}_{H_{1}}\left[\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}\leq c_{m,n}\right]$
$\leq\mathbb{P}_{H_{1}}\left[\left\|\rho-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}-\left\|\rho-\hat{\rho}\right\|_{1}\leq c_{m,n}\right]$
$=\mathbb{P}_{H_{1}}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}+\left\|\rho-\hat{\rho}\right\|_{1}\geq\left\|\rho-\sigma\right\|_{1}-c_{m,n}\right]$
$\leq\mathbb{P}_{H_{1}}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{2}\right]$
$\ \ +\mathbb{P}_{H_{1}}\left[\left\|\rho-\hat{\rho}\right\|_{1}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{2}\right].$
If $\left\|\rho-\sigma\right\|_{1}=c>c_{m,n}$, then for large enough $k$ we can continue the chain with
$\leq d\exp\left\{-\frac{m(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{344d^{3}}\right\}$
$\ +d\exp\left\{-\frac{n\left(\left\|\rho-\sigma\right\|_{1}-c_{m,n}\right)^{2}}{344d^{3}}\right\}$
$\leq 2d\exp\left\{-\frac{k\left(\left\|\rho-\sigma\right\|_{1}-c_{m,n}\right)^{2}}{344d^{3}}\right\}.$
Looking at the exponential term, we get
$\exp\left\{-\frac{k(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{344d^{3}}\right\}$
$=\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344d^{3}}+O\left(m^{1/2}\log(1/\alpha)\right)\right\}.$
Therefore
$\beta_{\rho,\sigma}(M_{m})$
$\leq\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344d^{3}}+O\left(k^{1/2}\log(1/\alpha)\right)\right\}.f$
A-H Proof of Theorem 6
Here, we will make use of Equation 30, to construct $\hat{\sigma}$.
Let our decision rule be
$\hat{H}_{0}:F(\rho,\hat{\sigma})$ $>1-c_{m}$
$\hat{H}_{1}:F(\rho,\hat{\sigma})$ $\leq 1-c_{m},$
where $c_{m}$ will again be a term used to control our type I error.
Now, we have
$\alpha(M_{m})$ $=\mathbb{P}\left[F(\rho,\hat{\sigma})\leq 1-c_{m}|\sigma=\rho\right]$
$\leq\exp\left\{-2mc_{m}+O(\log(m))\right\}.$
So, taking $c_{m}=\Theta\left(\frac{\log(m)}{m}\right)$, we get $\alpha(M_{m})\leq\alpha$.
On the other hand, under $H_{1}$ we will assume that $\left\|\rho-\sigma\right\|_{1}>0$, and we will denote probabilities conditioned on this event with $\mathbb{P}_{H_{1}}\left[\cdot\right]$. Then
$\beta_{\rho,\sigma}(M_{m})$ $=\mathbb{P}_{H_{1}}\left[F(\rho,\hat{\sigma})>1-c_{m}\right].$
Now, by Lemma 4 we have $F(\rho,\hat{\sigma})\leq 1-\frac{1}{4}\left\|\rho-\hat{\sigma}\right\|_{1}^{2}$, so that
$\mathbb{P}_{H_{1}}\left[F(\rho,\hat{\sigma})>1-c_{m}\right]$
$\leq\mathbb{P}_{H_{1}}\left[1-\frac{1}{4}\left\|\rho-\hat{\sigma}\right\|_{1}^{2}\geq 1-c_{m}\right]$
$=\mathbb{P}_{H_{1}}\left[\left\|\rho-\hat{\sigma}\right\|_{1}\leq\sqrt{4c_{m}}\right]$
$\leq\mathbb{P}_{H_{1}}\left[\left\|\rho-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}\leq\sqrt{4c_{m}}\right]$
$=\mathbb{P}_{H_{1}}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}\geq\left\|\rho-\sigma\right\|_{1}-\sqrt{4c_{m}}\right]$
$\leq\mathbb{P}_{H_{1}}\left[2(1-F(\sigma,\hat{\sigma}))^{1/2}\geq\left\|\rho-\sigma\right\|_{1}-\sqrt{4c_{m}}\right]$
$=\mathbb{P}_{H_{1}}\left[1-F(\sigma,\hat{\sigma})\geq\frac{\left(\left\|\rho-\sigma\right\|_{1}-\sqrt{4c_{m}}\right)^{2}}{4}\right],$
where in the last equality we require that $m$ is large enough such that $\left\|\rho-\sigma\right\|_{1}-\sqrt{4c_{m}}>0$. Then
$\implies\mathbb{P}_{H_{1}}\left[F(\rho,\hat{\sigma})>1-c_{m}\right]$
$\leq\exp\left\{-\frac{m\left(\left\|\rho-\sigma\right\|_{1}-\sqrt{4c_{m}}\right)^{2}}{2}+O(\log(m))\right\}$
$\leq\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}+O\left(m^{1/2}\log^{1/2}(m)\right)\right\}.$
A-I Proof of Theorem 7
Here, we will make use of Equation 30 to construct $\hat{\rho},\hat{\sigma}$ with guarantee from Equation 30.
Now, our decision rule will be
$\hat{H}_{0}$ $:\left\|\hat{\sigma}-\hat{\rho}\right\|_{1}<c_{m,n}$
$\hat{H}_{1}$ $:\left\|\hat{\sigma}-\hat{\rho}\right\|_{1}\geq c_{m,n}.$
Under $H_{0}$, and by Lemma 4, we have
$\left\|\hat{\sigma}-\hat{\rho}\right\|_{1}$ $\leq\left\|\hat{\sigma}-\sigma\right\|_{1}+\left\|\sigma-\hat{\rho}\right\|_{1}$
$=\left\|\hat{\sigma}-\sigma\right\|_{1}+\left\|\rho-\hat{\rho}\right\|_{1}$
$\leq 2\left(\sqrt{1-F(\hat{\sigma},\sigma)}+\sqrt{1-F(\hat{\rho},\rho)}\right),$
and
$\left(\sqrt{1-F(\hat{\rho},\rho)}+\sqrt{1-F(\hat{\sigma},\sigma)}\right)^{2}$
$\qquad\leq 3(1-F(\hat{\rho},\sigma))+3(1-F(\hat{\sigma},\sigma)),$
so that
$\mathbb{P}\left[\left\|\hat{\sigma}-\hat{\rho}\right\|_{1}\geq c_{m,n}\right]$
$=\mathbb{P}\left[\left\|\hat{\sigma}-\hat{\rho}\right\|_{1}^{2}\geq c_{m,n}^{2}\right]$
$\leq\mathbb{P}\left[4\left(\sqrt{1-F(\hat{\sigma},\sigma)}+\sqrt{1-F(\hat{\rho},\sigma)}\right)^{2}\geq c_{m,n}^{2}\right]$
[…]
$\leq\mathbb{P}\left[1-F(\hat{\sigma},\sigma)\geq\frac{c_{m,n}^{2}}{24}\right]+\mathbb{P}\left[1-F(\hat{\rho},\rho)\geq\frac{c_{m,n}^{2}}{24}\right]$
$\leq\exp\left\{-\frac{mc_{m,n}^{2}}{12}+O(d^{2}\log(m))\right\}$
$\ \ +\exp\left\{-\frac{nc_{m,n}^{2}}{12}+O(d^{2}\log(m))\right\}$
$\leq 2\exp\left\{-\frac{kc_{m,n}^{2}}{12}+O(d^{2}\log(k))\right\}.$
Taking $c_{m,n}=\Theta\left(\frac{\log(k)}{k^{1/2}}\right)$ gives us $\alpha(M_{m})\leq\alpha$.
Now (as before), under $H_{1}$, we assume that $\left\|\sigma-\rho\right\|_{1}=c>0$.
$\beta_{\rho,\sigma}(M_{m})$ $=\mathbb{P}_{H_{1}}\left[\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}\leq c_{m,n}\right],$
and
$\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}$
$\geq\left\|\hat{\rho}-\sigma\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}$
$\geq\left\|\rho-\sigma\right\|_{1}-\left\|\rho-\hat{\rho}\right\|_{1}-\left\|\sigma-\hat{\sigma}\right\|_{1}$
$\geq\left\|\rho-\sigma\right\|_{1}-2(1-F(\rho,\hat{\rho}))^{1/2}-2(1-F(\sigma,\hat{\sigma}))^{1/2};$
therefore,
$\mathbb{P}_{H_{1}}\left[\left\|\hat{\rho}-\hat{\sigma}\right\|_{1}\leq c_{m,n}\right]$
$\leq\mathbb{P}_{H_{1}}\left[\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{2}\leq(1-F(\rho,\hat{\rho}))^{1/2}+(1-F(\sigma,\hat{\sigma}))^{1/2}\right]$
$\leq\mathbb{P}_{H_{1}}\left[(1-F(\rho,\hat{\rho}))^{1/2}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{4}\right]$
$\ \ +\mathbb{P}_{H_{1}}\left[(1-F(\sigma,\hat{\sigma}))^{1/2}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{m,n}}{4}\right]$
$=\mathbb{P}_{H_{1}}\left[1-F(\rho,\hat{\rho})\geq\frac{(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{16}\right]$
$\ \ +\mathbb{P}_{H_{1}}\left[1-F(\sigma,\hat{\sigma})\geq\frac{(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{16}\right].$
Note that
$\left(\left\|\rho-\sigma\right\|_{1}-c_{m,n}\right)^{2}$ $=\left\|\rho-\sigma\right\|_{1}^{2}-2c_{m,n}\left\|\rho-\sigma\right\|_{1}-c_{m,n}^{2}$
$\geq\left\|\rho-\sigma\right\|_{1}-2c_{m,n},$
so that
$\mathbb{P}_{H_{1}}\left[1-F(\rho,\hat{\rho})\geq\frac{(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{16}\right]$
$\leq\exp\left\{-\frac{m(\left\|\rho-\sigma\right\|_{1}-c_{m,n})^{2}}{16}+O\left(\log(m)\right)\right\}$
$\leq\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}+O(k^{1/2}\log^{2}(k)\log(1/\alpha)\right\}.$
Hence
$\beta_{\rho,\sigma}(M_{m})$
$\leq\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}}{8}+O\left(k^{1/2}\log^{2}(k)\log(1/\alpha)\right)\right\}.$
A-J Sanov’s Theorem Result
By [17, Lemma 1], there exists a sequence of POVMs $\{A_{\rho,r_{m}},I-A_{\rho,r_{m}}\}_{m}$ such that
$\operatorname{Tr}[\rho^{\otimes m}(I-A_{\rho,r_{m}})]\leq(m+1)^{\frac{(d+4)(d-1)}{2}}e^{-m\cdot r_{m}}.$
Describing the POVM using notation in the paper, we would have
$A_{\rho,r_{m}}=\sum_{(p^{\prime},\rho^{\prime})\in S^{c}_{\rho,r_{m}}\cap\mathcal{R}_{m}[B]}T_{np^{\prime},\rho^{\prime},B}^{n}.$
Then we define $r_{m}$ so that the type I error is controlled, i.e.
$r_{m}$ $=\frac{(d+4)(d-1)}{2}\frac{\ln(m+1)}{m}+\frac{\ln(1/\alpha)}{m}$
$=O\left(\frac{\ln(m)}{m}\right),$
so that $\operatorname{Tr}[\rho^{\otimes m}(I-A_{\rho,r_{m}})]\leq\alpha$.
Now, if we assume that $\sigma\neq\rho$, then using [17, Corollary 6], for $0<R<\hat{D}(\rho\|\sigma)$, there exists $m$ large enough ($r_{m}$ small enough) such that
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln\left(\operatorname{Tr}[\sigma^{\otimes m}A_{\rho,r_{m}}]\right)\right\}\leq-R.$
Then, if we take our decision rule to be the measurement associated with this POVM, it would yield the stated result.